RC meeting and Run Plan, Monday Sept. 25, 2023
Return to the RC Meeting Calendar
- 1 RC Meeting Information
- 2 Useful Links
- 3 Procedures
- 4 Status Update
RC Meeting Information
Time: Monday - Friday at 3:30pm
- Location: Counting house conference room
- ZoomGov Meeting Connection Information
Hall A call-in instructions:
- Call RC first.
- Techs on site 7-4:30 Mon-Th, 7-11AM Fr. RC will call Hall A work coordinator (Zach) during normal working hours. After hours, RC calls Tech on-call.
- After 9pm everything non-emergency needs to be deferred. What is emergency and what's not? Call RC. RC talks with Mark to decide.
Hall Access procedure:
- Access check list (binder on the counting room desk) needs to be filled out before making an access.
- Scan (take pictures with phone) and post it to the logbook.
- Any work inside the SBS, BB boundary need to have PS locked out -- requires two QEWs to throw the big switch. Exception: field measurements.
New instruction for shift crew when changing target:
- Beam restoration instructions have been updated. Printed version is in the counting room.
- When we change target, make sure the SOFs of Target BPMs are set to the correct values.
- SOFs for different targets: Needs an update after restoring beam tonight (Note: SOFs for "No Target" has been updated and they are same as the He3 target values)
PSS issue was fixed yesterday, and we get beam back to our hall during the late swing shift (after ~48 hrs). Beam checkout procedure was completed around 4 am this morning, and started production run on polarized He3 cell > 37% polarization.
However we were having issue of Bigbite magnet tripping frequently (almost more than 6 time within 24 hrs) from yesterday. https://logbooks.jlab.org/entry/4186990 We were letting the magnet to cool down for a while and turning back ON until this morning. Tech on Call (Andrew) accessed hall this morning to quickly check the bigbite, and he found water leakage from the power supply https://logbooks.jlab.org/entry/4187094. We asked Radcon to have full survey of hall, and went to controlled access to fix the water leakage https://logbooks.jlab.org/entry/4187101.
Ellen and Andrew worked in the hall about 4 hrs to fix the water leakage. Summary of their work https://logbooks.jlab.org/entry/4187236. They informed that Bigbite magnet is ready, so we asked MCC to go to beam permit around 3pm. [Ellen asked not to turn on the bigbite magnet if it trips again, and she will look more details tomorrow afternoon]
As accelerator has planned for maintenance tomorrow (plan for testing in BSY starting 3 am tomorrow and Linac downstairs maintenance during the day shift). If you want to work in the hall tomorrow, please let me know.
- Maintenance day (Tuesday 09/26/2023)
Events collected :
- He3 data during owl shift: 0.02 C (corrected charge)
- accumulated corrected charge on He3 till now: 61.932 C
Hall A Subsystems Status
- Polarization: 43.2% (3:08 pm) [measurement after 5:30 am this morning, as bigbite magnet was off]
Requires hall access
- Cooling jet hardware limit (Todd/Ellen)
- SBS GEMs made DAQ unstable (GEM group)
- Urgent: Need a Shift leader for OWL shift of tomorrow and the day after tomorrow
Run plan for the weekend
- Production run on polarized He3 target @35uA beam current (can take higher current?)
- Alternative if bigbite is not ready!!! (zero field measurement with sieve slit IN)
==> List of items in case He3 production running isn’t ready:
- Take data on reference cell at 10 atm H2 (portions that add up to 1 day)
- BigBite Sieve slit => only during day (Techs needed)
- with 8 carbon foils and take optics for 1 hr at 40 uA (Sean/Andrew - confirm time)
- 1 hour for single foil (magnets off) => Chance BBcal HV setting to appropriate values
- BigBite/single foil location reconstruction (~20 mins?) - BigBite + HCAL detector experts on call consulting analysis experts like Andrew/Sean (magnet on)